Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 106     Entry time: 24 Mar 2023, 21:46
Author: Lisa Randolph 
Group: HED 
Subject: HED shift summary 23.-24.3. 

8.2 keV, 1-2 bunches, 600uJ

23.3.

  • did timing runs and optimized the beamline alignment
  • Karabo was working very slow
  • continued with sample alignment
  • performed several xfel+laser shots
  • opened the chamber and replaced samples

24.3.

  • aligned xfel and Relax to TCC
  • sample alignment
  • taking shots
ELOG V3.1.4-7c3fd00